# A Practical Cross-Platform, Multi-Algorithm Study of Quantum Optimisation for Configurational Analysis of Materials

### v2

@@ 5 Results
Method | Post-Selection | AR Post-Selection | User Runtime (s) | QPU Time (s) | |
Brute Force | N/A | N/A | N/A | ||
@@
We now discuss the results for each method and how they scale for larger sizes of the problem in detail. In the following sub-sections we first focus on the results of quantum annealing (Section 5.1), and then the results of VQE (Section 5.2). Classical results are discussed throughout both sections.
5.1 Quantum Annealing Results
@@
In Table 1 it can be seen that the QPU time for quantum annealing only makes up a fraction of the user runtime. In Appendix E we dissect the user runtime and show that the majority of this is coming from the embedding procedure (Figure 19(a)). This motivates us to explore more efficient embedding techniques, which we do in the next section as well as solving larger sizes of QUBO.
5.1.1 Scaling to larger problem sizes: Quantum Annealing
We now discuss how quantum annealing runtime scales as a function of problem size. The time spent by quantum annealing on each problem instance is partly determined by the number of anneals (shots) and the annealing time. The number of anneals is chosen so that our results are statistically significant, and the annealing time is found through hyperparameter optimisation. Figure 5 displays the user runtime for simulated annealing and quantum annealing (with two different embedding techniques) as a function of the number of QUBO variables, . The quantum annealing embedding technique used for the results in the previous section was minor embedding, D-Wave’s default mapping scheme, which is a heuristic method. An alternative embedding technique, ‘clique embedding’ [59], can be used for QUBO problems where the QUBO graph forms a ‘clique’, which means that it has full connectivity (100% density). Clique embedding is an algorithm that runs in polynomial time in the worst case[60, 61, 59], which has uniform chain lengths (or close to uniform) [60]. We observe near-constant scaling with clique embedding in Figure 19(a). At larger problem sizes this may start to grow polynomially.
As discussed previously, the primary contributor to the user runtime of quantum annealing using minor embedding is the embedding time, detailed in Appendix E and explicitly shown in Figure 19(a). As problem size increases, the required average qubit chain length also increases, shown in Figure 19(c). Consequently, the heuristic minor embedding technique must explore more mapping possibilities, leading to poor scaling. This can also explain the large standard deviation error bars observed in Figure 5: there is more variability possible when the minor heuristic embeds larger problems and requires longer qubit chains. A new embedding is found each time the minor heuristic is performed, which can give variable performance in terms of time and quality of solution. In future work, the minor embeddings which yield good performance could be saved and re-used for multiple experiments.
@@ 5.2 VQE Results
Strategies to reduce this user runtime could include using a less-optimised transpilation search at the expense of solution quality. Another recent technique introduced by IBM is fractional gates [71]. Fractional gates are essentially new native gates which remove the need for and rotation gates to be decomposed into many native gates. This can reduce circuit depth significantly. Our initial experimentation revealed that they can reduce the user runtime by up to a factor of two. Another technique that can be used to reduce gate depth is AI transpilation [72]. As discussed in Section 4, less strict convergence criteria would result in less iterations (and shorter user runtime), again sacrificing solution quality.
@@
5.2.1 Scaling to larger problem sizes: VQE QPU
Performing a scaling analysis for VQE on a real QPU proved challenging, primarily due to runtime increases as the problem size is scaled up. Beyond 50 variables, VQE failed to return non-zero . However, at the 72 variable problem size, convergence criteria were met and sub-optimal solutions were found (distribution and convergence plots are shown in Figures 13, 14, 15, 16 and 17, as well as the performance metrics for the different problem sizes in Table 4). At large problem sizes, more qubits are used and errors accumulate causing poor convergence to a state which has no overlap with the ground state. The extended runtime relative to the allocated device access time also meant that multiple experimental runs were not possible. Moreover, proper hyperparameter optimisation was not possible either due to this runtime, meaning that hyperparameters either need to be imported from smaller problem instances ran on the state vector solver or extrapolated from them.
@@ 5.2.1 Scaling to larger problem sizes: VQE QPU
Regarding near-term algorithms, it remains to either find structured instances of problems where provable speed-ups exist or to experimentally find possible modest speed-ups. It is also worth noting that using a one-qubit-per-variable encoding makes it impossible to scale to large problems with current hardware due to qubit count limitations. Efforts in multi-variable to qubit encoding are investigated in Refs. [75, 76]. In Appendix D.1 we discuss barren plateaus and classical simulability in the context of our ansatz.
6 Conclusions and Outlook
We introduced a systematic approach for analysing the performance of various algorithms for solving QUBO problems. This approach is applicable to other use-cases and problems in general. We applied it to the configurational analysis problem, determining the energy of defective graphene structures. We defined performance metrics, considering both solution quality and runtime, that are hardware and algorithm agnostic. We then explored the scaling of the problem to larger sizes. Our results show that simulated annealing was the best-performing method in both solution quality and runtime, with a polynomial scaling in time. It can also access large problems of hundreds of variables, in contrast to the quantum methods. We also find evidence of polynomial scaling with quantum annealing, solving QUBOs up to variables. Of the two embedding techniques that we tested, clique embedding was found to be optimal in terms of runtime.
We encountered challenges in scaling the VQE methods for several reasons, namely lack of convergence within the maximum number of iterations, convergence to states with no overlap with the ground state and limited experimental repetitions. However, we were able to solve a dense QUBO with 50 variables and obtain sub-optimal solutions for a 72 variable dense QUBO. A myriad of techniques that could further improve the performance of gate-based VQAs exist, such as:
@@ 6 Conclusions and Outlook
Quantum hardware is constantly improving, with many companies producing ambitious roadmaps promising more and high-quality physical qubits, and early steps towards quantum error correction. We leave finding potential speed-ups to future experimental works and hope that they can follow our framework for analysing the performance of different methods when solving optimisation problems.
@@ Acknowledgements
Phalgun Lolur conceived and planned the project. Kieran McDowall (KM) carried out the investigation. KM and Theodoros Kapourniotis analysed and interpreted the results with support from Chris Oliver. Konstantinos Georgopoulos was responsible for overall supervision of the work. KM prepared the manuscript which was then proof-read and revised by all authors.
@@ Appendix B Hyperparameter Search
Our hyperparameter search finds the set of parameters for each algorithm which return the highest probability of sampling the ground state. The hyperparameter optimisation is a grid-based search, conducted by defining feasible search spaces and checking different combinations of values. Other, more rigorous strategies for hyperparameter optimisation exist, such as random search and Bayesian optimisation. Use of these methods could yield better hyperparameters. Results are plotted for each algorithm in Figures 6, 7, 8, 9. Hyperparameters for VQE on the QPU were informed by VQE on the state vector solver. Other schemes exist to find the optimal penalty terms. A more systematic approach is explored in Ref. [99] for finding these terms.
@@ Appendix C Technical Information and Additional Data for Simulated Annealing
Method | Post-Selection | AR Post-Selection | User Runtime (s) | QPU Time (s) | |
Brute Force | N/A | N/A | N/A | ||
@@
Method | Post-Selection | AR Post-Selection | User Runtime (s) | QPU Time (s) | |
@@ Appendix D Technical Information and Additional Data for VQE
QUBO Variables | Post-Selection | User Runtime (minutes) | QPU Time (minutes) | |
The number of shots was set to 10,000 for both the real device and for the state vector solver. To solve the QUBO problem using VQE, each variable is initially mapped to a single physical qubit. Qiskit’s circuit transpilation is used when running on the real device, with the optimization_level=3 [101]. The qubit connectivity plot for ibm_fez is shown in Figure 10(a), where the qubits that were used are highlighted in red. The qubits that were used were chosen based off the error rates of each qubit, which can vary with device calibration. for VQE on the QPU should be minimal as qiskit serverless was used.
@@
The circuit ansatz used was Qiskit’s ‘RealAmplitudes’ [102]. For this study, only one alternating layer (repetition) was used, along with the default entanglement pattern ‘reverse_linear’. Measurements are made in the computational basis (single qubit Pauli-Z basis), such that a 0 or 1 is measured depending on the overlap with the Pauli-Z eigenstates and .
[…]
@C=1em @R=1em
[…]
D.1 Ansatz, Lie Algebra, Barren Plateaus and Classical Simulability
The ansatz used was qiskit’s RealAmplitudes circuit, shown in Figure 11, which is commonly used for chemistry applications or classification circuits in machine learning [102]. It is particularly suited for problems with real-valued solutions (since the ansatz only generates real amplitudes, no complex amplitudes). The circuit consists of alternating layers of Y rotations and CX entanglements. Figure 12 shows the ansatz for the 2 qubits.
[…]
The Dynamical Lie Algebra (DLA) for this circuit is given by all possible nested commutators between the generators , until no new independent skew-Hermitian operator is generated. The first two elements of the DLA are and (which we will write as and ), since . By noting the third element of g is . To find other elements of the DLA we now take the commutator of our current elements. We can do this in Pennylane [103] to give the DLA:
Considering the additional two gates produces no new elements in the DLA. Therefore . The dimension of the special unitary group is , which for gives . The Hilbert space is and for full universality, we require the available gates to span all of . That means when we have all unitaries of available to us, we can reach any state in Hilbert space from any other state. Therefore, the RealAmplitudes ansatz is not universal.
The dimension of the DLA for the RealAmplitudes with reverse linear entanglement scales polynomially with n. This can be seen by the fact that each contributes a local generator , therefore, n generators. Each gate introduces a new generator (e.g. ) on the target qubit. The number of these generators scales linearly with n because the entanglement is limited to neighbouring qubits. The new number of generators produced under commutation is restricted by the limited connectivity.
Ref. [104] states that barren plateaus do not exist for QUBO problems due to having 2-local observables. A key idea in Ref. [105] is that if there is an absence of barren plateaus this implies that your circuit can be simulated classically in polynomial time (rather than exponential time using the state vector simulator). The simulation algorithm: G-sim [106], exploits low dimensional lie algebra. In summary, our ansatz likely avoids barren plateaus and is efficiently classical simuable.
D.2 RealAmplitudes vs QAOA Ansatz (Hamiltonian agnostic vs informed)
@@
When comparing different ansatz in our hyperparameter search we found that the RealAmplitudes ansatz returned higher values of (results in Figure 8(e)). The simple explanation for this is that the RealAmplitudes ansatz has more parameters and is therefore more expressive allowing it to find a state with good overlap with the ground state. But what happens if we increase the depth of QAOA so that it has the same number of parameters as the RealAmplitudes ansatz? At , for the variable problem size, QAOA has just independent parameters, but already has a depth of gates, whereas the RealAmplitudes depth is gates (these are the abstract circuits before transpilation). This is due to the fact that our problem Hamiltonian , is fully connected. Consequently, the QAOA circuit has full connectivity (a fully connected entanglement arrangement), implementation of which on the IBM QPU is possible only with the introduction of SWAP gates. Running and simulating these QAOA circuits at can already take a long amount of time which makes increasing quite infeasible.
@@ Appendix E Technical Information and Additional Data for Quantum Annealing
To run a QUBO problem on the quantum annealer, each variable must be encoded to the hardware. Given the limited connectivity of the hardware, multiple physical qubits need to be mapped to a single variable. The mapping procedures used were minor embedding, D-Wave’s default mapping scheme, and clique embedding, intended for fully connected problem graphs. The minor embedding mapping to the D-Wave Advantage_system 6.4 QPU is shown in Figure 10(b). The time taken for this embedding procedure on the 18 variable problem was found to be . The percentage of broken chains for minor embedding on the 18 variable QUBO was %.
Figure 19(a) shows the embedding time required for all problem sizes solved. Figure 19(b) displays the QPU access time that the D-Wave device allocates for solving problem sizes of up to 72 variables, and Figure 19(c) presents the average chain lengths used. Full results for quantum annealing using both minor and clique embedding are shown in Table 5 for the larger problem sizes.
QUBO Variables | Embedding | Post-Selection | Runtime (s) | QPU Time (s) | |
Minor | |||||
Minor | |||||
Minor | |||||
Minor | |||||
Minor | |||||
Minor | |||||
Clique | |||||
Clique | |||||
Clique | |||||
Clique | |||||
Clique | |||||
Clique | |||||
Clique |
@@
Minor | ||||
Minor | ||||
Minor | ||||
Minor | ||||
Minor | ||||
Minor | ||||
Clique | ||||
Clique | ||||
Clique | ||||
Clique | ||||
Clique | ||||
Clique | ||||
Clique |

### v4

@@ V Results
| Method | Post-Sel. | AR Post-Sel. | User Runtime (s) | QPU Time (s) | |
| Brute Force | 1 | N/A | N/A | N/A | |
@@
We now discuss the results for each method and how they scale for larger sizes of the problem in detail. In the following sub-sections we first focus on the results of quantum annealing (Section V.1), and then the results of VQE (Section V.2). Classical results are discussed throughout both sections.
V.1 Quantum Annealing Results
@@
In Table 1 it can be seen that the QPU time for quantum annealing only makes up a fraction of the user runtime. In Appendix E we dissect the user runtime and show that the majority of this is coming from the embedding procedure (Figure 18a). This motivates us to explore more efficient embedding techniques, which we do in the next section as well as solving larger sizes of QUBO.
V.1.1 Scaling to larger problem sizes: Quantum Annealing
We now discuss how quantum annealing runtime scales as a function of problem size. The time spent by quantum annealing on each problem instance is partly determined by the number of anneals (shots) and the annealing time. The number of anneals is chosen so that our results are statistically significant, and the annealing time is found through hyperparameter optimisation. Figure 5 displays the user runtime for simulated annealing and quantum annealing (with two different embedding techniques) as a function of the number of QUBO variables, . The quantum annealing embedding technique used for the results in the previous section was minor embedding, D-Wave’s default mapping scheme, which is a heuristic method. An alternative embedding technique, ‘clique embedding’ [103], can be used for QUBO problems where the QUBO graph forms a ‘clique’, which means that it has full connectivity (100% density). Clique embedding is an algorithm that runs in polynomial time in the worst case[16, 66, 103], which has uniform chain lengths (or close to uniform) [16]. We observe near-constant scaling with clique embedding in Figure 18a. At larger problem sizes this may start to grow polynomially.
As discussed previously, the primary contributor to the user runtime of quantum annealing using minor embedding is the embedding time, detailed in Appendix E and explicitly shown in Figure 18a. As problem size increases, the required average qubit chain length also increases, shown in Figure 18c. Consequently, the heuristic minor embedding technique must explore more mapping possibilities, leading to poor scaling. This can also explain the large standard deviation error bars observed in Figure 5: there is more variability possible when the minor heuristic embeds larger problems and requires longer qubit chains. A new embedding is found each time the minor heuristic is performed, which can give variable performance in terms of time and quality of solution. In future work, the minor embeddings which yield good performance could be saved and re-used for multiple experiments.
@@ V.2 VQE Results
Strategies to reduce this user runtime could include using a less-optimised transpilation search at the expense of solution quality. Another recent technique introduced by IBM is fractional gates [79]. Fractional gates are essentially new native gates which remove the need for and rotation gates to be decomposed into many native gates. This can reduce circuit depth significantly. Our initial experimentation revealed that they can reduce the user runtime by up to a factor of two. Another technique that can be used to reduce gate depth is AI transpilation [34]. As discussed in Section IV, less strict convergence criteria would result in less iterations (and shorter user runtime), again sacrificing solution quality.
@@
V.2.1 Scaling to larger problem sizes: VQE QPU
Performing a scaling analysis for VQE on a real QPU proved challenging, primarily due to runtime increases as the problem size is scaled up. Beyond 50 variables, VQE failed to return non-zero . However, at the 72 variable problem size, convergence criteria were met and sub-optimal solutions were found (distribution and convergence plots are shown in Figures 12, 13, 14, 15 and 16, as well as the performance metrics for the different problem sizes in Table 4). At large problem sizes, more qubits are required and errors accumulate, causing poor convergence to a state which has no overlap with the ground state. The extended runtime relative to the allocated device access time also meant that multiple experimental runs were not possible. Moreover, thorough hyperparameter optimisation was not possible on the QPU due to these runtime constraints. QPU hyperparameters were chosen based on small problems ran on the state vector solver or extrapolated from those results.
@@ V.2.1 Scaling to larger problem sizes: VQE QPU
Regarding near-term algorithms, it remains to either find structured instances of problems where provable speed-ups exist or to experimentally find possible modest speed-ups. It is also worth noting that using a one-qubit-per-variable encoding makes it impossible to scale to large problems with current hardware due to qubit count limitations. Efforts in multi-variable to qubit encoding are investigated in Refs. [50, 29].
VI Conclusions and Outlook
We introduced a fair and systematic benchmarking approach for analysing the performance of various algorithms for solving QUBO problems. This approach is applicable to other use-cases and problems in general. We applied it to the configurational analysis problem, determining the energy of defective graphene structures. We defined performance metrics, considering both solution quality and runtime, that are hardware and algorithm agnostic. We then explored the scaling of the problem to larger sizes.
Our results show that simulated annealing was the best-performing method in both solution quality and runtime, exhibiting polynomial time scaling. It can also access large problems of hundreds of variables, in contrast to the quantum methods. We also find some provisional evidence of polynomial scaling with quantum annealing, solving QUBOs up to variables. Of the two embedding techniques that we tested, clique embedding was found to be optimal in terms of runtime due to its suitability for fully connected problems. We found that using small penalty coefficients, , which keep the QUBO energy range small, in combination with our post-selection, which removes any solutions that violate the constraint (due to using a small ) can return higher probability of obtaining the optimal solution for all methods.
We encountered challenges in scaling the VQE method for several reasons, namely lack of convergence within the maximum number of iterations, convergence to states with no overlap with the ground state and limited experimental repetitions due to large IBM runtimes. However, we were able to solve a dense QUBO with 50 variables and obtain sub-optimal solutions for a 72 variable dense QUBO. Using CVaR helped return a higher probability of obtaining optimal solutions with VQE.
Other techniques that could further improve the performance of gate-based VQAs exist, such as:
@@ VI Conclusions and Outlook
Classical surrogate training, in which variational parameters for the QPU are transferred over from those found in approximate classical simulation [47].
@@
Quantum hardware is constantly improving, with many companies producing ambitious roadmaps promising more high-quality physical qubits, and early steps towards quantum error correction. We leave finding potential speed-ups to future experimental works and hope that they can follow our framework for cross-platform benchmarking of different algorithms when solving optimisation problems.
@@ Acknowledgements
PL conceived, planned and supervised the project. KM carried out the investigation. KM and TK analysed and interpreted the results with support from CO. KG was responsible for overall supervision of the work. KM prepared the manuscript which was then proof-read and revised by all authors.
@@ Appendix B Hyperparameter Search
Our hyperparameter search finds the set of parameters for each algorithm which return the highest probability of sampling the ground state. The hyperparameter optimisation is a grid-based search, conducted by defining feasible search spaces and checking different combinations of values. Other, more rigorous strategies for hyperparameter optimisation exist, such as random search [89] and Bayesian optimisation. Use of these methods could yield better hyperparameters. Results are plotted for each algorithm in Figures 6, 7, 8, 9. Hyperparameters for VQE on the QPU were informed by VQE on the state vector solver. Other schemes exist to find the optimal penalty terms. A more systematic approach is explored in Ref. [35] for finding these terms.
@@ Appendix C Technical Information and Additional Data for Simulated Annealing
| Method | Post-Sel. | AR Post-Sel. | User Runtime (s) | QPU Time (s) | |
| Brute Force | 1 | N/A | N/A | N/A | |
@@
| Method | Post-Sel. | AR Post-Sel. | User Runtime (s) | QPU Time (s) | |
@@ Appendix D Technical Information and Additional Data for VQE
| QUBO Variables | Post-Selection | User Runtime (min) | QPU Time (min) | |
| 8 | 0.509 | 0.762 | 63.2 | 17.3 |
| 18 | ||||
| 32 | 0.0075 | 0.038 | 151 | 31.3 |
| 50 | 0.0007 | 0.016 | 137 | 42.4 |
| 72 | 0 | 0 | 252 | 98.2 |
The number of shots was set to 10,000 for both the real device and for the state vector solver. To solve the QUBO problem using VQE, each variable is initially mapped to a single physical qubit. Qiskit’s circuit transpilation is used when running on the real device, with the optimization_level=3 [78]. The qubit connectivity plot for ibm_fez is shown in Figure 10a, where the qubits that were used are highlighted in red. The qubits that were used were chosen based off the error rates of each qubit, which can vary with device calibration. for VQE on the QPU should be minimal as qiskit serverless was used.
@@
The circuit ansatz used was Qiskit’s ‘RealAmplitudes’ [99], shown in Figure 11, which is commonly used for chemistry applications or classification circuits in machine learning. It is particularly suited for problems with real-valued solutions (since the ansatz only generates real amplitudes, no complex amplitudes). The circuit consists of alternating layers of Y rotations and CX entanglements. For this study, only one alternating layer (repetition) was used, along with the default entanglement pattern ‘reverse_linear’. Measurements are made in the computational basis (single qubit Pauli-Z basis), such that a 0 or 1 is measured depending on the overlap with the Pauli-Z eigenstates and . For each problem size solved with VQE the distribution, post-selected distribution and convergence plots are shown in Figures 12, 13, 14, 15 and 16.
D.1 RealAmplitudes vs QAOA Ansatz (Hamiltonian agnostic vs informed)
@@
When comparing different ansatz in our hyperparameter search we found that the RealAmplitudes ansatz returned higher values of (results in Figure 8e). The simple explanation for this is that the RealAmplitudes ansatz has more parameters and is therefore more expressive allowing it to find a state with good overlap with the ground state. But what happens if we increase the depth of QAOA so that it has the same number of parameters as the RealAmplitudes ansatz? At , for the variable problem size, QAOA has just independent parameters, but already has a depth of gates, whereas the RealAmplitudes depth is gates (these are the abstract circuits before transpilation). This is due to the fact that our problem Hamiltonian , is fully connected. Consequently, the QAOA circuit has full connectivity (a fully connected entanglement arrangement), implementation of which on the IBM QPU is possible only with the introduction of SWAP gates. Running and simulating these QAOA circuits at can already take a long amount of time which makes increasing quite infeasible.
@@ Appendix E Technical Information and Additional Data for Quantum Annealing
To run a QUBO problem on the quantum annealer, each variable must be encoded to the hardware. Given the limited connectivity of the hardware, multiple physical qubits need to be mapped to a single variable. The mapping procedures used were minor embedding, D-Wave’s default mapping scheme, and clique embedding, intended for fully connected problem graphs. The minor embedding mapping to the D-Wave Advantage_system 6.4 QPU is shown in Figure 10b. The time taken for this embedding procedure on the 18 variable problem was found to be . The percentage of broken chains for minor embedding on the 18 variable QUBO was %.
Figure 18a shows the embedding time required for all problem sizes solved. Figure 18b displays the QPU access time that the D-Wave device allocates for solving problem sizes of up to 72 variables, and Figure 18c presents the average chain lengths used. Full results for quantum annealing using both minor and clique embedding are shown in Table 5 for the larger problem sizes.
| QUBO Variables | Embedding | Post-Sel. | Runtime (s) | QPU Time (s) | |
| 8 | Minor | ||||
| 18 | Minor | ||||
| 32 | Minor | ||||
| 50 | Minor | ||||
| 72 | Minor | ||||
| 98 | Minor | ||||
| 8 | Clique | ||||
| 18 | Clique | ||||
| 32 | Clique | ||||
| 50 | Clique | ||||
| 72 | Clique | ||||
| 98 | Clique | ||||
| 128 | Clique |
@@
| 8 | Minor | 3 | 4 | 600 |
| 18 | Minor | 1 | 3 | 1400 |
| 32 | Minor | 1 | 4 | 1600 |
| 50 | Minor | 1 | 6 | 1000 |
| 72 | Minor | 1 | 1 | 2000 |
| 98 | Minor | 1 | 1 | 2000 |
| 8 | Clique | 3 | 4 | 1000 |
| 18 | Clique | 1 | 3 | 2000 |
| 32 | Clique | 1 | 5 | 600 |
| 50 | Clique | 1 | 8 | 2000 |
| 72 | Clique | 1 | 13 | 2000 |
| 98 | Clique | 1 | 15 | 2000 |
| 128 | Clique | 1 | 22 | 2000 |